### v2

# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

Abstract.
In this paper, we develop a unified approach to study the intersection Betti numbers of moduli spaces of one-dimensional semistable sheaves on smooth projective surfaces. Assuming the irreducibility of such moduli spaces, we prove that their intersection Betti numbers in a certain range of degrees coincide with the stable Betti numbers of Hilbert schemes of points. As an application, for minimal surfaces of Kodaira dimension 0, we show that these intersection Betti numbers stabilize in each fixed degree, which fits into the broader context of stable cohomology for moduli spaces of sheaves. In the case of Enriques surfaces, we also prove a refined stabilization result related to Oberdieck’s conjecture on perverse Hodge numbers.
Key words and phrases:
moduli spaces of sheaves, intersection cohomology, perverse filtrations, compactified Jacobians
2020 Mathematics Subject Classification:
14D20, 14F06 (primary), 32S60 (secondary)
1. Introduction
In recent years, moduli spaces of one-dimensional sheaves on surfaces have attracted considerable attention. On the one hand, their connections with enumerative geometry and the theory of Higgs bundles have led to a number of new conjectures and results (e.g., [3, 4, 14, 20, 21, 25, 27, 35, 36, 46]). On the other hand, these moduli spaces exhibit not only asymptotic behavior analogous to that of moduli spaces of torsion-free sheaves, but also pathologies in contrast to the torsion-free case (see [42, §5]), revealing rich geometry that remains to be understood.
Let $S$ be a connected, smooth, complex projective surface with a fixed ample divisor $H$. For a nonzero effective divisor $\beta$ on $S$ and an integer $\chi$, let $M_{\beta,\chi}$ be the coarse moduli space parametrizing S-equivalence classes of one-dimensional semistable (with respect to $H$) sheaves $\mathcal{F}$} on $S$ with fixed determinant $\det({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and Euler characteristic $\chi({\mathcal{F}})=\chi$. The moduli space $M_{\beta,\chi}$ admits a Hilbert–Chow morphism
$h:M_{\beta,\chi}\to|\beta|$
to the linear system $|\beta|$, defined by taking the Fitting support. Over the open subscheme $|\beta|^{\mathrm{int}}\subset|\beta|$ parametrizing integral curves, the fibers of $h$ are compactified Jacobians.
A natural step toward understanding the geometry of $M_{\beta,\chi}$ is to determine its topological invariants. Since $M_{\beta,\chi}$ is singular in general, we consider its intersection cohomology, which is better-behaved than the ordinary cohomology. In this paper, we study the intersection Betti numbers of $M_{\beta,\chi}$, defined as the dimensions of the intersection cohomology groups $I\!H^{k}(M_{\beta,\chi})$ (with $\mathbb{Q}$}-coefficients) for $k\in\mathbb{Z}_{\geq 0}$. Our first main result expresses these intersection Betti numbers, in a certain range of degrees, as the coefficients $b_{k}^{\infty}$ of the formal power series
(1.1) $\sum_{k\geq 0}b_{k}^{\infty}q^{k}=\prod_{m\geq 1}\frac{(1+q^{2m-1})^{b_{1}(S)}(1+q^{2m+1})^{b_{1}(S)}}{(1-q^{2m})^{b_{2}(S)+1}(1-q^{2m+2})},$
where $b_{i}(S)$ denotes the $i$-th Betti number of $S$.
Theorem 1.1.
Suppose that $M_{\beta,\chi}$ is irreducible, and denote by
$c\vcentcolon=\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}}),M_{\beta,\chi})$
the codimension of $M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}})$ in $M_{\beta,\chi}$. If
  1. (1)
    $\beta$ is $\max\{1,2k-2\}$-very ample (see Definition 2.4), and
  2. (2)
    $k\leq\min\left\{c-1,\frac{2}{3}\dim|\beta|\right\}$,
then
$\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})=b_{k}^{\infty}.$
By Göttsche’s formula [11], the number $b_{k}^{\infty}$ is exactly the $k$-th Betti number $b_{k}(S^{[n]})$ of the Hilbert scheme $S^{[n]}$ of $n$ points on $S$ for any integer $n\geq k$. The connection between (intersection, virtual, or ordinary) Betti numbers of moduli spaces of sheaves on $S$ and $b_{k}^{\infty}$ has been observed in various settings. For some moduli spaces of torsion-free sheaves, this connection has been established, for instance, in [8, 23]. Similar results have been obtained for moduli spaces of one-dimensional sheaves on specific surfaces such as K3 surfaces [29], the projective plane [46], and del Pezzo surfaces [36]. The above Theorem 1.1 provides a unifying perspective on the intersection Betti numbers of moduli spaces of one-dimensional sheaves on surfaces.
To prove Theorem 1.1, we adapt the approach in Sections 2 and 3 of our previous joint work with Pi and Shen [36] to a more general setting. For del Pezzo (as well as K3 and Abelian) surfaces, the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ of $M_{\beta,\chi}$ is smooth. The main difficulty in generalizing the method in [36] to other cases arises from the singularities in $h^{-1}(|\beta|^{\mathrm{int}})$. Unlike in the torsion-free case, dimension bounds for the singular locus of $M_{\beta,\chi}$ were largely unknown, since standard techniques for torsion-free sheaves, such as those in [16, Ch. 9], are generally inapplicable to torsion sheaves. We use the deformation theory of curve singularities to show that the dimension of the singular locus of $h^{-1}(|\beta|^{\mathrm{int}})$ is governed by the positivity of $\beta$, which may be of independent interest and will be essential to our proof of Theorem 1.1. The precise statement is as follows.
Theorem 1.2.
If $\beta$ is $\max\{1,2k-2\}$-very ample, then there exists an open subscheme $U_{k}\subset|\beta|^{\mathrm{int}}$ such that
  1. (1)
    $\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1$;
  2. (2)
    the open subscheme $h^{-1}(U_{k})$ of $M_{\beta,\chi}$ is smooth.
Motivated by Theorem 1.1 and the results for torsion-free sheaves, we propose the following conjecture on the stabilization of intersection Betti numbers, which can be viewed as a rank 0 analogue of [8, Conjecture 1.1].
Conjecture 1.3.
Given any ample divisor $\beta_{0}$ on $S$ and any $k\in\mathbb{Z}_{\geq 0}$, $\chi\in\mathbb{Z}$, there exists $d(\beta_{0},k,\chi)\in\mathbb{Z}_{>0}$ such that for every integer $d\geq d(\beta_{0},k,\chi)$, the intersection Betti number $\dim_{\mathbb{Q}}I\!H^{k}(M_{d\beta_{0},\chi})$ is independent of $d$. If furthermore $M_{d\beta_{0},\chi}$ is irreducible, then the stable intersection Betti number is given by
(1.2) $\dim_{\mathbb{Q}}I\!H^{k}(M_{d\beta_{0},\chi})=b_{k}^{\infty}.$
As an application of Theorem 1.1, we prove that Conjecture 1.3 holds for all minimal surfaces of Kodaira dimension 0, namely, K3, Abelian, Enriques, and bielliptic surfaces. We focus primarily on the Enriques and bielliptic cases.
Theorem 1.4.
Suppose that $S$ is a minimal surface of Kodaira dimension 0. Then Conjecture 1.3 is true. Moreover, $d(\beta_{0},k,\chi)$ can be chosen independently of $\chi$ such that (1.2) holds for all integers $d\geq d(\beta_{0},k)\vcentcolon=d(\beta_{0},k,\chi)$.
Remark 1.5.
When $S$ is a K3 or an Abelian surface, the moduli space $M_{\beta,1}$ is smooth, and the statement of Conjecture 1.3 for $\chi=1$ can be deduced from known results in the literature, such as those in [45]. In the K3 case, Mauri and Migliorini [29] proved the stabilization of intersection Betti and intersection Hodge numbers of possibly singular $M_{\beta,\chi}$. The Abelian case can be proved similarly using the argument in [29, §10.3]. As in [29, 36], an analysis of $\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$ is needed for our proof of Theorem 1.4. However, our approach to determining the stable intersection Betti numbers is different from that in [29] and, in particular, does not involve the deformation equivalence of moduli spaces with Hilbert schemes of points on $S$. Such equivalence fails in the Enriques case (since in this case $M_{\beta,\chi}$ is odd-dimensional, e.g., [37]) and is unknown in the bielliptic case.
Although moduli spaces of sheaves on K3 and Abelian surfaces have been extensively studied, much less is known about the cases of Enriques and bielliptic surfaces. In the bielliptic case, some progress was made in [34].
When $S$ is a generic Enriques surface, assuming that $\beta^{2}>0$ and the class of $\beta$ is not divisible by $2$ in the group $\operatorname{Num}(S)$ of numerical equivalence classes of divisors on $S$, Saccà [37] proved that $M_{\beta}\vcentcolon=M_{\beta,1}$ is a smooth projective Calabi–Yau variety of dimension $\beta^{2}+1$ and computed its first few Betti numbers. In this case, although the odd cohomology of $M_{\beta}$ can be nonzero as shown in [37, Theorem 5.10], our Theorem 1.4 implies that for fixed odd $k$, the $k$-th Betti number $b_{k}(M_{\beta})$ vanishes if $\beta$ is sufficiently positive. We also prove, in this setting, a refined version (Theorem 1.6 below) of Theorem 1.4 on the stabilization of the invariants $n_{\beta}^{i,j}$ (defined by (2.6) with $\varphi$ taken to be $h:M_{\beta}\to|\beta|$). These invariants are conjectured to refine the Gromov–Witten/Pandharipande–Thomas invariants of the local surface $\operatorname{Tot}(K_{S})$ [15, 18, 27] and are the shifted perverse Hodge numbers considered in [35].
Theorem 1.6.
Suppose that $S$ is a generic Enriques surface. Given any ample divisor $\beta_{0}$ on $S$ with $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$ and any $i,j\in\mathbb{Z}_{\geq 0}$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ such that for all odd integers $d\geq d(\beta_{0},i,j)$,
$n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j},$
where $n_{\infty}^{i,j}$ is defined by the following identity of formal power series
(1.3) $\sum_{i,j\geq 0}n_{\infty}^{i,j}q^{i}t^{j}=(1-qt)\prod_{m\geq 1}\frac{1}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{10}(1-q^{m-1}t^{m+1})}.$
Remark 1.7.
Under the assumption of [35, Conjecture A], Oberdieck proved a stronger stabilization result [35, Proposition 1.1] for the shifted perverse Hodge numbers. We note that the product formula (1.3) coincides with the one in [35, Proposition 1.1]. It remains open whether our approach, which does not rely on [35, Conjecture A], can be modified to show $n_{\beta}^{i,j}=n^{i,j}_{\infty}$ if $i<\beta^{2}/4$ and $j<\beta^{2}/4-1$. An explicit expression for $d(\beta_{0},i,j)$ in Theorem 1.6 will be given in (5.24).
The rest of this paper is organized as follows. In Section 2, we collect some notation, definitions, and results needed later. Section 3 is devoted to the local versality of the universal family of curves in high codimension and contains the proof of Theorem 1.2. In Section 4, we prove Theorem 1.1 and discuss some difficulties arising in the general case. Finally, in Section 5, we study in detail the cases of minimal surfaces of Kodaira dimension 0 to prove Theorems 1.4 and 1.6, concluding with a discussion of related open questions.
Conventions
All schemes are assumed to be of finite type over the field $\mathbb{C}$ of complex numbers. A variety is an integral and separated scheme. A curve is a projective scheme of pure dimension $1$. A surface is a variety of dimension $2$. By a point of a scheme, we always mean a closed point. All sheaves are assumed to be coherent. For a sheaf $\mathcal{E}$} on a projective scheme $X$, we denote by $h^{i}({\mathcal{E}})$ (resp. $\operatorname{ext}^{i}({\mathcal{E}},{\mathcal{E}})$) the dimension of the $i$-th sheaf cohomology $H^{i}(X,{\mathcal{E}})$ (resp. $i$-th Ext group $\operatorname{Ext}^{i}({\mathcal{E}},{\mathcal{E}})$) as a $\mathbb{C}$-vector space.
2. Preliminaries
2.1. Moduli spaces of one-dimensional sheaves
We first review some basic notions concerning the moduli spaces studied in this paper and then discuss some properties regarding the Hilbert–Chow morphism and irreducibility.
Fix a smooth polarized surface $(S,H)$ as in Section 1. Recall that the dimension of a sheaf is the dimension of its support. A one-dimensional sheaf $\mathcal{F}$} on $S$ is called semistable (resp. stable) (with respect to $H$) if every nonzero subsheaf $\mathcal{G}$$\subsetneq{\mathcal{F}}$ is one-dimensional with
$\frac{\chi({\mathcal{G}})}{c_{1}({\mathcal{G}})\cdot H}\leq(\text{resp.}<)\,\frac{\chi({\mathcal{F}})}{c_{1}({\mathcal{F}})\cdot H}.$
Two semistable sheaves are called S-equivalent if their Jordan–Hölder filtrations have isomorphic factors up to reordering. Each S-equivalence class is represented by a unique polystable sheaf, i.e., a semistable sheaf which is the direct sum of stable sheaves. In general, the coarse moduli space $M_{\beta,\chi}$ considered in this paper depends on the ample divisor $H$, but we have omitted $H$ from the notation since $H$ is fixed throughout.
Let ${\mathcal{M}}_{\beta,\chi}$ be the moduli stack of one-dimensional semistable sheaves on $S$ with $\det({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and $\chi({\mathcal{F}})=\chi$. Choose a large integer $n$ such that for any semistable sheaf $\mathcal{F}$} in ${\mathcal{M}}_{\beta,\chi}$, the sheaf ${\mathcal{F}}(nH)$ is globally generated and $h^{j}({\mathcal{F}}(nH))=0$ for $j>0$. Set $N\vcentcolon=(\beta\cdot H)n+\chi$. Denote by $Q$ the Quot scheme parametrizing quotients ${\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{E}}$, where $\mathcal{E}$} is semistable with fixed Hilbert polynomial $\chi({\mathcal{E}}(mH))=(\beta\cdot H)m+\chi$ and the induced map $H^{0}(S,{\mathcal{O}}_{S}^{\oplus N})\to H^{0}(S,{\mathcal{E}}(nH))$ is an isomorphism.
Let $\operatorname{Pic}(S)$ denote the Picard scheme of $S$. There is a morphism
$\det:Q\to\operatorname{Pic}(S)$
sending ${\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{E}}]$ to $\det({\mathcal{E}})$. Then ${\mathcal{M}}_{\beta,\chi}$ has a quotient stack presentation
${\mathcal{M}}_{\beta,\chi}\simeq[Q_{\beta}/\operatorname{GL}(N)],$
where $Q_{\beta}\vcentcolon=\det^{-1}({\mathcal{O}}_{S}(\beta))$ and $\operatorname{GL}(N)$ is the general linear group acting naturally on $Q_{\beta}$.
The natural morphism
$\nu:{\mathcal{M}}_{\beta,\chi}\to M_{\beta,\chi}$
is the good moduli space for ${\mathcal{M}}_{\beta,\chi}$. Denote by
$\tilde{h}=h\circ\nu:{\mathcal{M}}_{\beta,\chi}\to|\beta|$
the stacky Hilbert–Chow morphism. Since the fibers of $h$ over the locus $|\beta|^{\mathrm{int}}$ of integral curves are compactified Jacobians and have dimension $p_{a}(\beta)$, where $p_{a}(\beta)$ is the arithmetic genus of $\beta$, the fibers of $\tilde{h}$ over $|\beta|^{\mathrm{int}}$ have dimension $p_{a}(\beta)-1$. Whether the fiber dimension of $\tilde{h}$ jumps depends on numerical properties of the canonical divisor $K_{S}$.
Lemma 2.1 ([46, Corollary 1.3]).
If $K_{S}$ is nef and $|\beta|^{\mathrm{int}}\neq\varnothing$, then for any curve $C\in|\beta|$, the fiber $\tilde{h}^{-1}(C)$ has dimension $p_{a}(\beta)-1$.
This property of $\tilde{h}$ gives constraints on the geometry of $M_{\beta,\chi}$. The next proposition generalizes [25, Theorem 2.3] and [46, Theorem 1.5] beyond the del Pezzo case.
Proposition 2.2.
Suppose that $K_{S}$ is nef and $|\beta|$ contains a smooth connected curve. If $\dim|\beta|=\chi({\mathcal{O}}_{S}(\beta))-1$, then
  1. (i)
    the moduli space $M_{\beta,\chi}$ is irreducible;
  2. (ii)
    all fibers of $h:M_{\beta,\chi}\to|\beta|$ have dimension $p_{a}(\beta)$.
Proof.
(i) It suffices to show that $Q_{\beta}$ is irreducible. We use the deformation theory of sheaves and quotients (see [9, Theorem 6.4.9, Proposition 6.5.1] and [16, 2.A.8, Theorem 4.5.3]). For any $\xi\in Q_{\beta}$ representing the short exact sequence
(2.4) $0\to{\mathcal{K}}\to{\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{F}}\to 0,$
the dimension of $Q_{\beta}$ at $\xi$, denoted by $\dim_{\xi}Q_{\beta}$, satisfies (cf. [16, Proposition 4.5.9])
(2.5) $\dim_{\xi}Q_{\beta}$ $\geq\dim_{\xi}Q-\dim\operatorname{Pic}(S)$
$\geq\operatorname{ext}^{0}({\mathcal{K}},{\mathcal{F}})-(\operatorname{ext}^{2}({\mathcal{F}},{\mathcal{F}})-h^{2}({\mathcal{O}}_{S}))-h^{1}({\mathcal{O}}_{S})$
$=N^{2}+\beta^{2}+\chi({\mathcal{O}}_{S})-1,$
where the last equality follows from the long exact sequence of Ext groups for (2.4) and the Riemann–Roch formula. By [42, Proposition 2.2], the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ is irreducible and contains the smooth open subscheme $h^{-1}(|\beta|^{\mathrm{sm}})$, where $|\beta|^{\mathrm{sm}}\subset|\beta|$ is the open subscheme parametrizing smooth curves. Then by [16, Corollary 4.3.5], the open subscheme $Q_{\beta}^{\mathrm{int}}\subset Q_{\beta}$ parametrizing those quotient sheaves with integral support is irreducible and contains the smooth open subscheme $Q_{\beta}^{\mathrm{sm}}$ parametrizing those quotient sheaves with smooth support. There is a unique irreducible component $Y\subset Q_{\beta}$ containing $Q_{\beta}^{\mathrm{int}}$. If $Y\not=Q_{\beta}$, by taking an irreducible component $Y^{\prime}\neq Y$ and noting that $Y^{\prime}\cap Q_{\beta}^{\mathrm{sm}}=\varnothing$, we deduce from Lemma 2.1 and [43, Tag 0DS4] that
$\dim Y^{\prime}<\dim\operatorname{GL}(N)+\dim|\beta|+(p_{a}(\beta)-1)=N^{2}+\beta^{2}+\chi({\mathcal{O}}_{S})-1,$
which contradicts (2.5). Hence, $Q_{\beta}=Y$ is irreducible, and so is $M_{\beta,\chi}$.
(ii) By the upper semicontinuity of the fiber dimension of $h$, $\dim h^{-1}(C)\geq p_{a}(\beta)$ for any curve $C\in|\beta|$. Note that ${\mathcal{M}}_{\beta,\chi}$ is also irreducible since $Q_{\beta}$ is. Consider the morphism $\nu:{\mathcal{M}}_{\beta,\chi}\to M_{\beta,\chi}$ and its restriction $\nu|_{\tilde{h}^{-1}(C)}:\tilde{h}^{-1}(C)\to h^{-1}(C)$. By Lemma 2.1 and [43, Tag 0DS4], we have
$\dim\tilde{h}^{-1}(C)-\dim h^{-1}(C)=(p_{a}(\beta)-1)-\dim h^{-1}(C)\geq-1=\dim{\mathcal{M}}_{\beta,\chi}-\dim M_{\beta,\chi},$
which forces $\dim h^{-1}(C)=p_{a}(\beta)$. ∎
Remark 2.3.
If $\beta$ is an ample divisor on a K3 (resp. an Abelian) surface $S$ satisfying $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)\geq 2$, Proposition 2.2 (ii) can also be obtained from the symplectic geometry of $M_{\beta,\chi}$: [^1] The authors are grateful to Mirko Mauri for pointing out this alternative approach and the relevant references via private communication. Using [5, Propositions 3.5, 3.14], one can show that $M_{\beta,\chi}$ (resp. $K_{\beta,\chi}$) is a primitive symplectic variety (e.g., [5, Definition 3.3]) admitting a Lagrangian fibration over $|\beta|$ (with no restrictions on $H$ or $\chi$), and thus all fibers of $h:M_{\beta,\chi}\to|\beta|$ have dimension $p_{a}(\beta)$ by [38, Theorem 3]. Here, when $S$ is an Abelian surface with identity $0_{S}$, we denote by $K_{\beta,\chi}$ the fiber $\widehat{\varphi}^{\,-1}(0_{S})$ of the isotrivial fibration $\widehat{\varphi}:M_{\beta,\chi}\to S$ (cf. [45, Definition 4.1]).
2.2. Intersection cohomology and perverse filtrations
For a quasi-projective variety $X$, let $\operatorname{IC}_{X}$ denote the intersection cohomology complex of $X$ with $\mathbb{Q}$}-coefficients and middle perversity. The intersection cohomology groups of $X$ are defined as the hypercohomology groups of $\operatorname{IC}_{X}$:
$I\!H^{k}(X)\vcentcolon={\mathbb{H}}^{k-\dim X}(X,\operatorname{IC}_{X}).$
Note that $I\!H^{k}(X)$ can also be defined if $X$ is allowed to be non-reduced (by the topological invariance of intersection cohomology) or reducible (even not pure dimensional, as shown in [6, §4.6]). When $X$ is smooth, $I\!H^{k}(X)$ can be identified with the rational cohomology group $H^{k}(X)$.
Let $\varphi:X\to Y$ be a proper morphism between quasi-projective varieties. We assume that $\dim X=a$, $\dim Y=b$, and all fibers of $\varphi$ have dimension aba-b. The perverse filtration
$P_{0}I\!H^{k}(X)\subset P_{1}I\!H^{k}(X)\subset\dots\subset I\!H^{k}(X)$
induced by $\varphi$ is an increasing filtration on $I\!H^{k}(X)$, governed by the topology of the morphism $\varphi$. It is defined by
$P_{i}I\!H^{k}(X)\vcentcolon=\mathrm{Im}\left\{{\mathbb{H}}^{k-b}(Y,{{}^{\mathbf{p}}\tau_{\leq i}}(R\varphi_{*}\operatorname{IC}_{X}[b-a]))\to{\mathbb{H}}^{k-b}(Y,R\varphi_{*}\operatorname{IC}_{X}[b-a])\right\},$
where $^{\mathbf{p}}\tau_{\leq\bullet}$ is the perverse truncation functor. We denote the dimension of the graded piece $\mathrm{Gr}_{i}^{P}I\!H^{i+j}(X)\vcentcolon=P_{i}I\!H^{i+j}(X)/P_{i-1}I\!H^{i+j}(X)$ by
(2.6) $n_{\varphi}^{i,j}\vcentcolon=\dim_{\mathbb{Q}}\mathrm{Gr}_{i}^{P}I\!H^{i+j}(X).$
When $X$ is smooth, the decomposition theorem [2] applied to $\varphi:X\to Y$ yields
$R\varphi_{*}{\mathbb{Q}}_{X}[b]\simeq\bigoplus_{i=0}^{2a-2b}\mathcal{P}_{i}[-i],$
where $\mathcal{P}_{i}$ are semisimple perverse sheaves on $Y$; in this case, $n_{\varphi}^{i,j}=\dim_{\mathbb{Q}}{\mathbb{H}}^{j-b}(Y,{\mathcal{P}}_{i})$.
For our purposes, we mainly consider the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$ and its restrictions. The invariants $n_{\varphi}^{i,j}$ with $\varphi$ taken to be (the restrictions of) $h$ will play a central role in Section 4.
2.3. Positivity conditions for divisors
To study the asymptotic behavior of $M_{\beta,\chi}$, we recall the definition of $k$-very ampleness for divisors on $S$, which generalizes base-point-freeness (0-very ampleness) and very ampleness ($1$-very ampleness).
Definition 2.4.
An effective divisor $\beta$ on $S$ is called $k$-very ample if for any 0-dimensional closed subscheme $Z\subset S$ of length $k+1$, the restriction map
$r_{Z}:H^{0}(S,{\mathcal{O}}_{S}(\beta))\to H^{0}(Z,{\mathcal{O}}_{S}(\beta)|_{Z})$
is surjective. We adopt the convention that every divisor is (1)(-1)-very ample.
When $h^{1}({\mathcal{O}}_{S}(\beta))=0$, the restriction map $r_{Z}$ is surjective if and only if $h^{1}({\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))=0$ by the long exact sequence of cohomology groups for the short exact sequence
$0\to{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta)\to{\mathcal{O}}_{S}(\beta)\to{\mathcal{O}}_{S}(\beta)|_{Z}\to 0,$
where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$. Definition 2.4 immediately implies two properties, which will be used in the subsequent sections:
  1. (1)
    If $\beta$ is $k$-very ample and $k^{\prime}\leq k$ $(k^{\prime}\in\mathbb{Z}_{\geq 0})$, then $\beta$ is also $k^{\prime}$-very ample.
  2. (2)
    If $\beta$ is ample, then for any $k\in\mathbb{Z}_{\geq 0}$, there exists $d_{k}\in\mathbb{Z}_{>0}$ such that $d\beta$ is $k$-very ample for every integer $d\geq d_{k}$.
3. Singular locus of relative compactified Jacobians
The goal of this section is to prove Theorem 1.2. We begin by reviewing some basic notions of plane curve singularities and deformation theory.
3.1. Singularities and deformations of locally planar curves
3.1.1. Singularity invariants
We first recall some analytic invariants of isolated singularities in locally planar curves.
Definition 3.1.
Let $f\in\mathbb{C}[[x,y]]$ be a convergent power series.
  1. (i)
    The numbers μ(f):=dim_CC[[x,y]]/(∂_x f,∂_y f), τ(f):=dim_CC[[x,y]]/(f,∂_x f,∂_y f) are called the Milnor number and Tjurina number of $f$, respectively.
  2. (ii)
    If $f$ is reduced, let O=C[[x,y]]/(f)↪O be the normalization of $\mathcal{O}$}. Then the number δ(f):=dim_CO/O is called the $\delta$-invariant of $f$.
The Milnor number and the $\delta$-invariant are related by the following formula of Milnor.
Proposition 3.2 ([32, §10]).
Let $f\in(x,y)\subset\mathbb{C}[[x,y]]$ be a reduced, convergent power series. Denote by $r(f)$ the number of irreducible factors of $f$. Then
(3.7) $\mu(f)=2\delta(f)-r(f)+1.$
If $C$ is a reduced curve on $S$, then the singular locus $C_{\mathrm{sing}}$ of $C$ is discrete. For any point $p\in C_{\mathrm{sing}}$, there exists a polynomial $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors such that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$, i.e.,
$\widehat{\mathcal{O}}_{C,p}\cong\mathbb{C}[[x,y]]/(f).$
Thus, the Milnor number, the Tjurina number, and the $\delta$-invariant of $C$ at $p$ can be defined as those of $f$ in Definition 3.1, denoted by $\mu_{p}(C)$, $\tau_{p}(C)$, and $\delta_{p}(C)$, respectively.
For completeness, we give a proof of the following standard interpretation (cf. [13, Ch. II, Exercise 1.4.2]) of the total Tjurina number of $C$ in terms of the Ext sheaf ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$, where $\Omega_{C}$ is the cotangent sheaf of $C$.
Lemma 3.3.
If $C$ is a reduced curve on S, then
$\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)=h^{0}({\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})).$
Proof.
The sheaf $\Omega_{C}$ is locally free over the smooth locus $C\setminus C_{\mathrm{sing}}$ of $C$, so ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$ is a torsion sheaf supported on $C_{\mathrm{sing}}$. Let $p\in C_{\mathrm{sing}}$ and assume that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$ for $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors. The lemma can be proved locally around $p$. For $R\vcentcolon=\mathbb{C}[x,y]/(f)$, we have $\Omega_{R}\cong(Rdx\oplus Rdy)/(\partial_{x}fdx+\partial_{y}fdy)$. Using the conormal exact sequence
$0\to(f)/(f^{2})\to R^{\oplus 2}\to\Omega_{R}\to 0,$
we obtain $\operatorname{Ext}^{1}(\Omega_{R},R)\cong R/(\partial_{x}f,\partial_{y}f)$, from which the result follows. ∎
3.1.2. Smoothness from local versality
Let $\pi_{B}:{\mathcal{C}}_{B}\to B$ be a family of reduced, locally planar curves, i.e., a proper and flat morphism whose fibers are reduced curves with at worst planar singularities. Assume that $B$ is a smooth variety. There is a natural map (cf.  [13, Ch. II, Lemma 1.20, Exercise 1.4.4])
(3.8) $\kappa_{\pi_{B},b}^{\text{loc}}:T_{b}B\to\prod_{p\in({\mathcal{C}}_{b})_{\mathrm{sing}}}T\mathrm{Def}_{{\mathcal{C}}_{b},p}=H^{0}({\mathcal{C}}_{b},{\mathcal{E}xt}^{1}(\Omega_{{\mathcal{C}}_{b}},{\mathcal{O}}_{{\mathcal{C}}_{b}})),$
called the local Kodaira–Spencer map, where $({\mathcal{C}}_{b})_{\mathrm{sing}}$ is the singular locus of ${\mathcal{C}}_{b}\vcentcolon=\pi_{B}^{-1}(b)$, and $T\mathrm{Def}_{{\mathcal{C}}_{b},p}$ is the tangent space of the deformation functor of the complete local $\mathbb{C}$-algebra $\widehat{{\mathcal{O}}}_{{\mathcal{C}}_{b},p}$.
Definition 3.4.
The family $\pi_{B}:{\mathcal{C}}_{B}\to B$ is called locally versal if for any point $b\in B$, the local Kodaira–Spencer map $\kappa_{\pi_{B},b}^{\text{loc}}$ in (3.8) is surjective.
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a family of integral, locally planar curves, then for any $e\in\mathbb{Z}$, denote by $\varrho_{B,e}:\overline{J}_{B}^{e}\to B$ the relative compactified Jacobian of degree $e$ associated with $\pi_{B}$. The fiber $\varrho_{B,e}^{-1}(b)$ over a point $b\in B$ is the moduli space of rank $1$, torsion-free sheaves of degree $e$ (or equivalently, with Euler characteristic $e+1-p_{a}({\mathcal{C}}_{b})$) on ${\mathcal{C}}_{b}$. Since $\pi_{B}$ admits, étale locally on $B$, a section which avoids the singularities of fibers of $\pi_{B}$, the smoothness of $\overline{J}_{B}^{e}$ is equivalent to that of $\overline{J}_{B}^{0}$. Hence, the following result follows directly from [10, Corollary B.3].
Proposition 3.5 ([10, Corollary B.3]).
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a locally versal family of integral, locally planar curves, then $\overline{J}_{B}^{e}$ is smooth for any $e\in\mathbb{Z}$.
3.2. Local versality and smoothness in high codimension
Now we show that the universal family $\pi:{\mathcal{C}}\to|\beta|$ of curves in $|\beta|$ is locally versal in sufficiently high codimension if $\beta$ is positive enough.
Proposition 3.6.
If $\beta$ is $\max\{1,2k-2\}$-very ample, then there exists an open subscheme $V_{k}\subset|\beta|$ such that
  1. (1)
    every point of $V_{k}$ corresponds to a reduced curve;
  2. (2)
    $\operatorname{codim}(|\beta|\setminus V_{k},|\beta|)\geq k+1$;
  3. (3)
    the universal family $\pi|_{\pi^{-1}(V_{k})}:\pi^{-1}(V_{k})\to V_{k}$ of curves in $V_{k}$ is locally versal.
Proof.
By setting $V_{0}\vcentcolon=V_{1}$, we may assume $k>0$. The assumption on $\beta$ implies that $\beta$ is both $k$-very ample and (2k22k-2)-very ample. We use an argument in [22] to control the locus of curves that are not $k$-nodal. Here, a curve is called $k$-nodal if it is reduced and has no singularities other than $k$ nodes. Since $\beta$ is $k$-very ample, by the proof of [22, Proposition 2.1], there is an open subscheme $V_{k}\subset|\beta|$ satisfying the following conditions:
  • $\operatorname{codim}(|\beta|\setminus V_{k},|\beta|)>k$.
  • Curves in $V_{k}$ are either $k$-nodal, or reduced of geometric genus $p_{a}(\beta)-k$.
Note that for a curve $C\in V_{k}$ which is not $k$-nodal, its geometric genus $p_{g}(C)$ satisfies
$p_{g}(C)=p_{a}(\beta)-\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)>p_{a}(\beta)-k.$
From this and Milnor’s formula (3.7), we obtain
(3.9) $\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)\leq\sum_{p\in C_{\mathrm{sing}}}\mu_{p}(C)\leq 2\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)<2k.$
It remains to show Property (3). For any curve $D\in V_{k}$, let $s_{D}\in H^{0}(S,{\mathcal{O}}_{S}(\beta))$ be a defining section of $D$. The map $\kappa_{\pi,D}^{\text{loc}}$ as in (3.8) can be identified with the map
$\bar{r}_{Z_{D}}:H^{0}(S,{\mathcal{O}}_{S}(\beta))/\langle s_{D}\rangle\to H^{0}(Z_{D},{\mathcal{O}}_{S}(\beta)|_{Z_{D}})$
induced by restriction to a 0-dimensional closed subscheme $Z_{D}\subset S$. By (3.9) (which still holds for a $k$-nodal curve) and Lemma 3.3, the length $h^{0}({\mathcal{O}}_{Z_{D}})$ of $Z_{D}$ satisfies
$h^{0}({\mathcal{O}}_{Z_{D}})=h^{0}({\mathcal{E}xt}^{1}(\Omega_{D},{\mathcal{O}}_{D}))=\sum_{p\in D_{\mathrm{sing}}}\tau_{p}(D)<2k.$
Hence, $\kappa_{\pi,D}^{\text{loc}}$ is surjective by the (2k22k-2)-very ampleness of $\beta$. ∎
Combining the above proposition with Proposition 3.5, we can prove Theorem 1.2.
Proof of Theorem 1.2.
Let $V_{k}$ be as in Proposition 3.6. Take $U_{k}=V_{k}\cap|\beta|^{\mathrm{int}}$. By Property (2) of $V_{k}$ in Proposition 3.6, we have
$\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1.$
It remains to show that $h^{-1}(U_{k})$ is smooth. Note that $h^{-1}(U_{k})$ is the relative compactified Jacobian of degree $p_{a}(\beta)-1+\chi$ associated with the universal family of curves in $U_{k}$. This universal family is locally versal by Property (3) of $V_{k}$ in Proposition 3.6. Thus, the smoothness of $h^{-1}(U_{k})$ follows from Proposition 3.5. ∎
4. Intersection Betti numbers
In this section, we generalize the approach in [36] to prove Theorem 1.1. To start with, we review some properties of relative Hilbert schemes of points over $|\beta|$, which will play an important role in the proof.
4.1. Relative Hilbert schemes of points
Let $\pi:{\mathcal{C}}\to|\beta|$ be the universal family of curves in $|\beta|$. For $k\in\mathbb{Z}_{\geq 0}$, denote by
$\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$
the relative Hilbert scheme of $k$ points on the fibers of $\pi$. For $C\in|\beta|$, the fiber of $\pi^{[k]}$ over $C$ is the Hilbert scheme $C^{[k]}$ parametrizing 0-dimensional, length $k$ closed subschemes of $C$. Note that $\pi^{[0]}:|\beta|\to|\beta|$ is the identity and $\pi=\pi^{[1]}:{\mathcal{C}}\to|\beta|$. The following property of $\pi^{[k]}$ is a corollary of [24, Theorem 1.1] (see also [42, Proposition 2.7]).
Lemma 4.1.
The morphism $\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$ has fibers of the same dimension $k$.
There is a natural morphism $\sigma_{k}:{\mathcal{C}}^{[k+1]}\to S^{[k+1]}$ defined by
$Z\subset C]\mapsto[Z\subset S],$
where $C\in|\beta|$, and $Z\subset C$ is a 0-dimensional closed subscheme of length $k+1$. The fiber of $\sigma_{k}$ over $Z\subset S]\in S^{[k+1]}$ is the projective space $\mathbb{P}(H^{0}(S,{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))^{\vee})$, where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$. A well-known connection between $k$-very ampleness and the structure of ${\mathcal{C}}^{[k+1]}$ as a scheme over $S^{[k+1]}$ via $\sigma_{k}$ is as follows (e.g., [42, Proposition 2.6]).
Lemma 4.2.
If $\beta$ is $k$-very ample, then ${\mathcal{C}}^{[k+1]}$ is a projective bundle over $S^{[k+1]}$.
Recall that for the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$, the fibers of $h$ over the locus $|\beta|^{\mathrm{int}}$ of integral curves are the compactified Jacobians of degree $p_{a}(\beta)-1+\chi$ and have dimension $p_{a}(\beta)$. For any open subscheme $U\subset|\beta|$, let $h_{U}:h^{-1}(U)\to U$ be the restriction of $h$, and write ${\mathcal{C}}_{U}^{[k]}$ for $(\pi^{[k]})^{-1}(U)$. Later, we will use the following result, due to Maulik–Yun [28, Theorem 2.13] and Migliorini–Shende [31, Corollary 2].
Theorem 4.3 ([28, 31] [^2] The results of [28, 31] apply since the smoothness of $h^{-1}(U)$ implies, via a section of the universal curve (étale locally on $U$, avoiding the singularities of fibers), that the relative compactified Jacobians $\overline{J}_{U}^{e}$ for all $e\in\mathbb{Z}$ are smooth, and by [41, Proposition 14], so are the relative Hilbert schemes ${\mathcal{C}}^{[k]}_{U}$ for all $k\in\mathbb{Z}_{\geq 0}$.).
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $P_{\bullet}H^{*}(h^{-1}(U))$ be the perverse filtration induced by $h_{U}:h^{-1}(U)\to U$. For any $k,\ell\in\mathbb{Z}_{\geq 0}$, we have an isomorphism of vector spaces
(4.10) $H^{\ell}\left({\mathcal{C}}^{[k]}_{U}\right)\cong\bigoplus_{i+j\leq k;\,i,j\geq 0}\mathrm{Gr}_{i}^{P}H^{\ell-2j}\left(h^{-1}(U)\right).$
Remark 4.4.
When $k=0$, the isomorphism (4.10) reads
$H^{\ell}(U)\cong\mathrm{Gr}_{0}^{P}H^{\ell}\left(h^{-1}(U)\right),$
which can be obtained as a consequence of Ngô’s support theorem [33, Théorème 7.2.1] (see [28, Theorem 2.4, §2.5] or [31, Corollary 9]). Ngô’s support theorem implies that $h_{U}:h^{-1}(U)\to U$ has full support. More precisely, there is an isomorphism
$Rh_{U*}{\mathbb{Q}}_{h^{-1}(U)}[\dim U]\simeq\bigoplus_{i=0}^{2p_{a}(\beta)}\operatorname{IC}(\wedge^{i}R^{1}\pi^{s}_{*}{\mathbb{Q}}_{{\mathcal{C}}_{U}^{\mathrm{sm}}})[-i],$
where $\pi^{s}:{\mathcal{C}}_{U}^{\mathrm{sm}}\to U^{\mathrm{sm}}$ is the universal family of smooth curves in $U$.
4.2. Strategy of the proof
For any open subscheme $U\subset|\beta|$, let
$N(U)\vcentcolon=2\operatorname{codim}(|\beta|\setminus U,|\beta|)-2.$
We simply write $N(\beta)$ for $N(|\beta|^{\mathrm{int}})$.
As a preliminary step, we focus on a smooth open subscheme $h^{-1}(U)$ of $M_{\beta,\chi}$ and prove the following result concerning graded dimensions of the perverse filtration induced by the restriction $h_{U}:h^{-1}(U)\to U$ of $h:M_{\beta,\chi}\to|\beta|$ (see (2.6)).
Theorem 4.5.
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth. For any $i,j\in\mathbb{Z}_{\geq 0}$ satisfying $i+j\leq N(U)$ and $3i+j\leq 2\dim|\beta|$, if $\beta$ is (i1)(i-1)-very ample, then
$n_{h_{U}}^{i,j}=[H(q,t)]^{i,j},$
where $H(q,t)]^{i,j}$ is the coefficient of $q^{i}t^{j}$ in the expansion of
(4.11) $H(q,t)\vcentcolon=(1-qt)\prod_{m\geq 1}\frac{(1+q^{m}t^{m-1})^{b_{1}(S)}(1+q^{m}t^{m+1})^{b_{1}(S)}}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{b_{2}(S)}(1-q^{m-1}t^{m+1})}.$
Remark 4.6.
Recall Göttsche’s formula [11] for $\sum_{k\geq 0}\sum_{i\geq 0}b_{i}(S^{[k]})z^{i}w^{k}$:
(4.12) $G(z,w)\vcentcolon=\prod_{m\geq 1}\frac{(1+z^{2m-1}w^{m})^{b_{1}(S)}(1+z^{2m+1}w^{m})^{b_{1}(S)}}{(1-z^{2m-2}w^{m})(1-z^{2m}w^{m})^{b_{2}(S)}(1-z^{2m+2}w^{m})}.$
After the change of variables
$z=t,\quad w=\frac{q}{t},$
the product formula $H(q,t)$ in (4.11) satisfies the following equality:
$\frac{H(q,t)}{1-qt}=G(z,w)\cdot\frac{1-w}{1-z^{2}}.$
By (2.6), the $k$-th Betti number of $h^{-1}(U)$ satisfies
(4.13) $b_{k}(h^{-1}(U))=\sum_{i=0}^{k}n_{h_{U}}^{i,k-i}.$
Since the right-hand side of (1.1) is exactly $H(q,q)$, the next corollary follows immediately from Theorem 4.5 and (4.13).
Corollary 4.7.
In the setting of Theorem 4.5, if $k\leq\min\{N(U),2\dim|\beta|/3\}$ and $\beta$ is (k1k-1)-very ample, then
$b_{k}(h^{-1}(U))=b_{k}^{\infty},$
where $b^{\infty}_{k}$ is defined by (1.1).
We prove Theorem 4.5 by relating $h^{-1}(U)$ to the relative Hilbert schemes ${\mathcal{C}}^{[k]}_{U}$. By (4.10), for any $k,\ell\in\mathbb{Z}_{\geq 0}$,
(4.14) $b_{\ell}({\mathcal{C}}^{[k]}_{U})=\sum_{i+j\leq k;\,i,j\geq 0}n_{h_{U}}^{i,\ell-i-2j}.$
To calculate the Betti numbers of ${\mathcal{C}}^{[k]}_{U}$, we use Lemmas 4.1 and 4.2.
Proposition 4.8.
In the setting of Theorem 4.5, if $\ell\leq\min\{N(U),2\dim|\beta|-2k\}$ and $\beta$ is (k1k-1)-very ample, then
(4.15) $b_{\ell}({\mathcal{C}}^{[k]}_{U})=\sum_{r=0}^{\lfloor\ell/2\rfloor}b_{\ell-2r}(S^{[k]}).$
Proof.
Since all fibers of $\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$ have dimension $k$ by Lemma 4.1, we have
$\operatorname{codim}({\mathcal{C}}^{[k]}\setminus{\mathcal{C}}_{U}^{[k]},{\mathcal{C}}^{[k]})=\operatorname{codim}(|\beta|\setminus U,|\beta|).$
By Lemma 4.2, the (k1)(k-1)-very ampleness of $\beta$ implies that ${\mathcal{C}}^{[k]}$ is a projective bundle over $S^{[k]}$. In particular, ${\mathcal{C}}^{[k]}$ is irreducible and smooth. Then it follows from [36, Lemma 2.4] (or the proof of Lemma 5.3) that for $\ell\leq N(U)$,
$b_{\ell}({\mathcal{C}}^{[k]}_{U})=b_{\ell}({\mathcal{C}}^{[k]}).$
Combining this with
$\dim{\mathcal{C}}^{[k]}-\dim S^{[k]}=\dim|\beta|-k\geq\ell/2,$
which follows from our assumption, we conclude that
$b_{\ell}({\mathcal{C}}^{[k]}_{U})=b_{\ell}({\mathcal{C}}^{[k]})=b_{\ell}(S^{[k]}\times\mathbb{P}^{\dim|\beta|-k})=\sum_{r=0}^{\lfloor\ell/2\rfloor}b_{\ell-2r}(S^{[k]}),$
thereby completing the proof of (4.15). ∎
Since all Betti numbers of $S^{[k]}$ are determined by Göttsche’s formula (4.12), now we can prove Theorem 4.5 by combining (4.14) and (4.15).
Proof of Theorem 4.5.
We use induction on $i$. For $i=0$ and $j\leq\min\{N(U),2\dim|\beta|\}$,
$n_{h_{U}}^{0,j}$ $=\dim_{\mathbb{Q}}\mathrm{Gr}_{0}^{P}H^{j}(h^{-1}(U))$
[…]
which proves the induction base. Then suppose that $n_{h_{U}}^{i^{\prime},j^{\prime}}=[H(q,t)]^{i^{\prime},j^{\prime}}$ holds for any $i^{\prime},j^{\prime}\in\mathbb{Z}_{\geq 0}$ satisfying $i^{\prime}<i$ and $i^{\prime}+j^{\prime}\leq i+j$. For any $\ell\leq i+j$, it follows from (4.14), our induction hypothesis, and (4.15) that
$n_{h_{U}}^{i,\ell-i}$ $=\sum_{i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
$=\left(b_{\ell}({\mathcal{C}}^{[i]}_{U})-b_{\ell}({\mathcal{C}}^{[i-1]}_{U})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}[H(q,t)]^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
$=\sum_{r=0}^{\lfloor\ell/2\rfloor}\left(b_{\ell-2r}(S^{[i]})-b_{\ell-2r}(S^{[i-1]})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}[H(q,t)]^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
$=[H(q,t)]^{i,\ell-i},$
where the last identity is essentially equivalent to [36, (31)] and follows from Remark 4.6. This completes the induction step and hence the proof. ∎
Finally, we prove our first main theorem stated in Section 1.
Proof of Theorem 1.1.
By the upper semicontinuity of fiber dimension, we have
(4.16) $N(\beta)\geq 2c-2.$
Since $\beta$ is $\max\{1,2k-2\}$-very ample, let $U_{k}\subset|\beta|^{\mathrm{int}}$ be as in Theorem 1.2. Using (4.16), the assumption $k\leq c-1$, and Theorem 1.2, we obtain the following properties:
  1. (1)
    $N(U_{k})\geq\min\{N(\beta),2k\}=2k$;
  2. (2)
    $h^{-1}(U_{k})$ is smooth.
Note that $|\beta|$ contains a connected smooth curve by Bertini’s theorem. Applying Corollary 4.7 with $U=U_{k}$, we have
(4.17) $b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty}.$
Since fibers of $h$ over integral curves have the same dimension, Theorem 1.2 implies
$\operatorname{codim}(h^{-1}(|\beta|^{\mathrm{int}})\setminus h^{-1}(U_{k}),h^{-1}(|\beta|^{\mathrm{int}}))=\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1.$
This, together with the assumption on $c=\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}}),M_{\beta,\chi})$, yields
$\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(U_{k}),M_{\beta,\chi})\geq\min\{c,k+1\}>k.$
Then by [30, Theorem 6.7.4] and the smoothness of $h^{-1}(U_{k})$, we have isomorphisms
$I\!H^{k}(M_{\beta,\chi})\cong I\!H^{k}(h^{-1}(U_{k}))\cong H^{k}(h^{-1}(U_{k})),$
and therefore, by (4.17),
$\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})=b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty},$
which completes the proof. ∎
4.3. Discussion on the general case
In general, the moduli space $M_{\beta,\chi}$ can be reducible even if $\beta$ is sufficiently positive (see [42, §5]). Moreover, even if we assume that $M_{\beta,\chi}$ is irreducible, it might not be true that the codimension $c$ in Theorem 1.1 is the same as $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$. In the cases where $c=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, the asymptotic behavior of $I\!H^{k}(M_{\beta,\chi})$ can be described more precisely, provided that $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, or equivalently $N(\beta)$ (defined in §4.2), can be calculated.
The following lemma allows us to estimate $N(\beta)$ by analyzing the decompositions of $\beta$.
Lemma 4.9.
Fix an ample divisor $\beta$ on $S$. There are finitely many pairs of classes $(\theta_{1},\theta_{2})$ in $\operatorname{Num}(S)$ such that $\theta_{i}$ ($i=1,2$) are the classes of nonzero effective divisors $C_{i}$ with
$C_{1}+C_{2}=\beta.$
Proof.
Let $D_{1},\cdots,D_{\rho}$ be an orthogonal basis for $\operatorname{Num}(S)_{\mathbb{Q}}\vcentcolon=\operatorname{Num}(S)\otimes_{\mathbb{Z}}{\mathbb{Q}}$, where $D_{1}$ is the class of an ample divisor. By the Hodge index theorem, $D_{\ell}^{2}<0$ for $1<\ell\leq\rho$. Given any $C_{1}$, $C_{2}$ as in the statement, write
$\beta\equiv\sum_{j=1}^{\rho}a_{j}D_{j},\quad C_{i}\equiv\sum_{j=1}^{\rho}a_{i,j}D_{j}$
for some $a_{j},a_{i,j}\in{\mathbb{Q}}$, where $\equiv$ denotes the numerical equivalence. For $1<\ell\leq\rho$, let $n_{\ell}>0$ be an integer such that $A_{\ell}\vcentcolon=n_{\ell}D_{1}+D_{\ell}$ is ample. Then $D_{1}\cdot C_{i}>0$ and $A_{\ell}\cdot C_{i}>0$, which implies
$a_{i,1}>0\quad\text{and}\quad n_{\ell}a_{i,1}D_{1}^{2}+a_{i,\ell}D_{\ell}^{2}>0.$
Combined with $D_{\ell}^{2}<0$ and $C_{1}+C_{2}=\beta$, this shows that the possible values of $a_{i,j}$ ($i=1,2$; $j=1,\cdots,\rho$) are bounded since $\beta$ is fixed, and thus the result follows. ∎
Motivated by the del Pezzo case ([36, Proposition 1.5]), along with the cases studied in Section 5, we ask the following question.
Question 4.10.
For any ample divisor $\beta$ on an arbitrary smooth projective surface $S$, is it true that
$\lim_{d\to+\infty}N(d\beta)=+\infty?$
5. Stabilization for minimal surfaces of Kodaira dimension zero
Throughout this section, except in §5.4, we consider a minimal surface $S$ of Kodaira dimension 0. In particular, its canonical divisor $K_{S}$ is numerically trivial. In this situation, the following proposition gives an affirmative answer to Question 4.10, which we then use to prove Theorem 1.4.
Proposition 5.1.
If $\beta$ is an ample divisor on a minimal surface $S$ of Kodaira dimension 0, then
$\lim_{d\to+\infty}N(d\beta)=+\infty.$
The proof of Proposition 5.1 will be given in §5.1.2, §5.2.2, and §5.3, where we analyze the Enriques, bielliptic, K3 and Abelian cases separately.
Together with Proposition 2.2 and Proposition 5.1, Theorem 1.1 implies the stabilization of the intersection Betti numbers of $M_{\beta,\chi}$.
Proof of Theorem 1.4.
Choose $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ such that for all $d\geq d(\beta_{0},k)$, the following properties hold:
  1. (1)
    $N(d\beta_{0})\geq 2k$ (guaranteed by Proposition 5.1);
  2. (2)
    $d\beta_{0}$ is $\max\{1,2k-2\}$-very ample;
  3. (3)
    $2\dim|d\beta_{0}|=2\chi({\mathcal{O}}_{S}(d\beta_{0}))-2\geq 3k$.
By Proposition 2.2 (i), $M_{d\beta_{0},\chi}$ is irreducible, and by Proposition 2.2 (ii), we have
$\operatorname{codim}(M_{d\beta_{0},\chi}\setminus h^{-1}(|d\beta_{0}|^{\mathrm{int}}),M_{d\beta_{0},\chi})$ $=\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$
$=1+\frac{1}{2}N(d\beta_{0})\geq k+1.$
Hence, the result follows from Theorem 1.1 with $\beta=d\beta_{0}$. ∎
5.1. The case of Enriques surfaces
5.1.1. Basic properties
Recall that an Enriques surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=h^{2}({\mathcal{O}}_{S})=0$. Its second Betti number is $b_{2}(S)=10$. The following property of linear systems on $S$ is well-known (e.g., [19, Theorem]) and was used in [37, Proposition 1.14] to compute the codimension of the locus of reducible curves in a linear system.
Lemma 5.2.
Let $D$ be a nonzero effective nef divisor on an Enriques surface $S$.
  1. (i)
    If $D^{2}>0$, then
    $\dim|D|=\chi({\mathcal{O}}_{S}(D))-1=\frac{D^{2}}{2}.$
  2. (ii)
    If $D^{2}=0$, then $D$ is linearly equivalent to either $mE$ or $mE+K_{S}$ for some $m\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E$ with $E^{2}=0$. Moreover,
    $\dim|D|=\begin{cases}\lfloor\frac{m}{2}\rfloor&(\text{if }D\sim mE)\\ \lfloor\frac{m-1}{2}\rfloor&(\text{if }D\sim mE+K_{S}).\end{cases}$
5.1.2. Proof of Proposition 5.1 in the Enriques case
By Lemma 4.9, there are finitely many pairs of nonzero effective divisors $(C_{1},C_{2})$ (up to linear or numerical equivalence) with $d\beta=C_{1}+C_{2}$. Any non-integral curve in $|d\beta|$ lies in the image of $|C_{1}|\times|C_{2}|$ for one of these pairs.
First we assume that $C_{1}$ and $C_{2}$ have nef irreducible components. To prove the lemma, we estimate all possible values of
$\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|.$
Case 1.1: $C_{1}^{2}>0$ and $C_{2}^{2}>0$. In this case, $C_{1}^{2}\geq 2$ and $C_{2}^{2}\geq 2$. By Lemma 5.2 (i) and the Hodge index theorem, we have
$\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $=C_{1}\cdot C_{2}\geq\sqrt{C_{1}^{2}C_{2}^{2}}$
$\geq\sqrt{2(d^{2}\beta^{2}-2C_{1}\cdot C_{2}-2)},$
where the last inequality follows from $d^{2}\beta^{2}=C_{1}^{2}+C_{2}^{2}+2C_{1}\cdot C_{2}$. Therefore,
(5.18) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)=C_{1}\cdot C_{2}\geq d\sqrt{2\beta^{2}}-2.$
Case 1.2: $C_{1}^{2}=0$ and $C_{2}^{2}>0$. By Lemma 5.2 (ii), $C_{1}$ is linearly equivalent to $m_{1}E_{1}$ or $m_{1}E_{1}+K_{S}$ for some $m_{1}\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E_{1}$ with $E_{1}^{2}=0$. Moreover,
(5.19) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{m_{1}}{2}\rfloor$
$\geq m_{1}\left(d\beta\cdot E_{1}-\frac{1}{2}\right)\geq d-\frac{1}{2}.$
Case 1.3: $C_{1}^{2}=C_{2}^{2}=0$. Let $m_{1}$ and $E_{1}$ be as in Case 1.2. By Lemma 5.2 (ii), $C_{2}$ is linearly equivalent to $m_{2}E_{2}$ or $m_{2}E_{2}+K_{S}$ for some $m_{2}\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E_{2}$ with $E_{2}^{2}=0$. Then we have $C_{1}\cdot C_{2}=m_{1}m_{2}E_{1}\cdot E_{2}=d^{2}\beta^{2}/2>0$ and
(5.20) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{m_{1}}{2}\rfloor-\lfloor\frac{m_{2}}{2}\rfloor\geq\frac{d^{2}\beta^{2}-(m_{1}+m_{2})}{2}$
$\geq\frac{d^{2}\beta^{2}-(1+d^{2}\beta^{2}/2)}{2}=\frac{d^{2}\beta^{2}-2}{4}.$
It remains to consider the case where every irreducible component of $C_{1}$ (resp. $C_{2}$) is nef (resp. a (2)(-2)-curve). By Lemma 4.9, for a fixed $d$, there are finitely many such $C_{2}$. In this case, it suffices to estimate
$\dim|d\beta|-\dim|C_{1}|.$
Case 2.1: $C_{1}^{2}>0$. By Lemma 5.2 (i),
(5.21) $\dim|d\beta|-\dim|C_{1}|=\frac{d\beta\cdot C_{2}+C_{1}\cdot C_{2}}{2}\geq\frac{d}{2}.$
Case 2.2: $C_{1}^{2}=0$. Let $m_{1}$ and $E_{1}$ be as in Case 1.2. By Lemma 5.2 (ii),
(5.22) $\dim|d\beta|-\dim|C_{1}|\geq\frac{d^{2}\beta^{2}}{2}-\lfloor\frac{m_{1}}{2}\rfloor\geq\frac{m_{1}(d\beta\cdot E_{1}-1)+d\beta\cdot C_{2}}{2}\geq d-\frac{1}{2}.$
Combining (5.18)–(5.22) concludes the proof. ∎
5.1.3. Refined invariants
Assume that $S$ is a generic Enriques surface. For an ample divisor $\beta$ whose class in $\operatorname{Num}(S)$ is not divisible by $2$, the moduli space $M_{\beta}=M_{\beta,1}$ is a smooth projective Calabi–Yau variety by [37]. The linear system $|\beta|$ contains a connected smooth curve (e.g., [37, Corollary 1.13]). Denote by $n_{\beta}^{i,j}$ the number defined by (2.6) with $\varphi$ taken to be $h:M_{\beta}\to|\beta|$.
In order to apply Theorem 4.5 to compute $n_{\beta}^{i,j}$, we need the following lemma which is parallel to [26, Corollary 5.3]. For the reader’s convenience, we provide a proof below.
Lemma 5.3.
Let $\beta$ be an ample divisor on a generic Enriques surface $S$ with $2\nmid\beta$ in $\operatorname{Num}(S)$. For any $i,j\in\mathbb{Z}_{\geq 0}$ satisfying $i+j\leq N(\beta)$, we have
$n_{\beta}^{i,j}=n^{i,j}_{h^{\circ}},$
where $h^{\circ}=h_{|\beta|^{\mathrm{int}}}:h^{-1}(|\beta|^{\mathrm{int}})\to|\beta|^{\mathrm{int}}$ is the restriction of $h:M_{\beta}\to|\beta|$.
Proof.
Let $W=M_{\beta}\setminus h^{-1}(|\beta|^{\mathrm{int}})$. Denote by $\iota:W\hookrightarrow M_{\beta}$ the closed embedding, and by $\jmath:h^{-1}(|\beta|^{\mathrm{int}})\hookrightarrow M_{\beta}$, $\jmath^{\circ}:|\beta|^{\mathrm{int}}\hookrightarrow|\beta|$ the open embeddings. By Proposition 2.2 (ii),
$\operatorname{codim}(W,M_{\beta})=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)=1+\frac{1}{2}N(\beta).$
Consider the exact triangle (all functors are derived)
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath_{*}\jmath^{*}{\mathbb{Q}}_{M_{\beta}}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}.$
By the property of Verdier duals [30, Proposition 5.3.9] and the smoothness of $M_{\beta}$,
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\simeq(D_{M_{\beta}}\iota_{*}{\mathbb{Q}}_{W})[-2\dim M_{\beta}],$
and therefore for $k\leq N(\beta)+1$,
${\mathbb{H}}^{k}(M_{\beta},\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})\cong H^{2\dim M_{\beta}-k}(W)^{\vee}=0.$
By the long exact sequence in hypercohomology, the vanishing of ${\mathbb{H}}^{k}(M_{\beta},\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})$ shows that for $\ell\leq N(\beta)$, the restriction map induces an isomorphism
(5.23) $H^{\ell}(M_{\beta})\cong H^{\ell}(h^{-1}(|\beta|^{\mathrm{int}})).$
Since the restriction to $h^{-1}(|\beta|^{\mathrm{int}})$ clearly preserves a relatively ample class (relative to $h$), it follows from [7, Lemma 3.3] that $n_{\beta}^{i,j}\leq n_{h^{\circ}}^{i,j}$ for $i+j\leq N(\beta)$. Then the result follows by (4.13) and (5.23). ∎
Now we can prove the stabilization of $n_{\beta}^{i,j}$ for moduli spaces of one-dimensional sheaves on a generic Enriques surface.
Proof of Theorem 1.6.
Choose $d(\beta_{0},i,j)$ such that for all $d\geq d(\beta_{0},i,j)$, the following properties hold:
  1. (1)
    $N(d\beta_{0})\geq i+j$ (guaranteed by Proposition 5.1);
  2. (2)
    $d\beta_{0}$ is (i1)(i-1)-very ample;
  3. (3)
    $d^{2}\beta_{0}^{2}\geq 3i+j$.
Property (3) is equivalent to $2\dim|d\beta_{0}|\geq 3i+j$ by Lemma 5.2 (i). Applying Lemma 5.3 and Theorem 4.5 with $\beta=d\beta_{0}$ and $U=|d\beta_{0}|^{\mathrm{int}}$, we have
$n_{d\beta_{0}}^{i,j}=n_{h_{U}}^{i,j}=[H(q,t)]^{i,j}=n_{\infty}^{i,j},$
which completes the proof. ∎
Again by (4.13), the next corollary follows immediately.
Corollary 5.4.
Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ with $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$ and any $k\in\mathbb{Z}_{\geq 0}$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ such that
$b_{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all $d\geq d(\beta_{0},k)$ with }2\nmid d,$
where $b_{k}^{\infty}$ is defined by (1.1).
Remark $5.$
Note that a generic Enriques surface does not contain any (2)(-2)-curves. By (5.18)–(5.20) and [44, Proposition 2.5], we can take
(5.24) $d(\beta_{0},i,j)=\max\left\{i+1,\,\left\lceil\frac{i+j+2}{2}\right\rceil,\,\left\lceil\frac{i+j+6}{2\sqrt{2\beta_{0}^{2}}}\right\rceil,\,\left\lceil\sqrt{\frac{2i+2j+6}{\beta_{0}^{2}}}\,\right\rceil\right\}$
in Theorem 1.6.
5.2. The case of bielliptic surfaces
5.2.1. Basic properties
Recall that a bielliptic surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=1$ and $h^{2}({\mathcal{O}}_{S})=0$. In this case, the exponential sequence reads
$0\to\operatorname{Pic}^{0}(S)\to\operatorname{Pic}(S)\to H^{2}(S,\mathbb{Z})\to 0,$
where $\operatorname{Pic}^{0}(S)=H^{1}(S,{\mathcal{O}}_{S})/H^{1}(S,\mathbb{Z})$ is the group of algebraically trivial divisors modulo linear equivalence and has the structure of an elliptic curve. By [1, pp. 112–114] (see also [39, Proposition 1.1]), there exist two elliptic curves $A$, $B$, together with a finite Abelian group $G$ acting on $A$ and $B$, such that
  1. (1)
    $A/G$ is an elliptic curve and $B/G\cong\mathbb{P}^{1}$;
  2. (2)
    $S\cong(A\times B)/G$, where $G$ acts on $A\times B$ componentwise.
We follow the convention in [39] to denote also by $A$ (resp. $B$) the class in $\operatorname{Num}(S)$ of a smooth fiber of the natural projection $\Psi:S\to B/G$ (resp. $\Phi:S\to A/G$). Then
$A^{2}=B^{2}=0,\quad A\cdot B=\gamma,$
where $\gamma$ is the order of $G$. There exist $\lambda,\mu\in{\mathbb{Q}}_{>0}$ such that the classes $\lambda A$ and $\mu B$ form a basis of $\operatorname{Num}(S)$ (see [39, Theorem 1.4] for the values of $\lambda$ and $\mu$).
We need the following lemma to compute the dimension of linear systems on $S$.
Lemma 5.6 ([39, Lemma 1.3]).
Let $D$ be a divisor of numerical class $uA+vB$ with $u,v\in{\mathbb{Q}}$ on a bielliptic surface $S$. The following properties hold:
  1. (i)
    $\chi(\mathcal{O}_{S}(D))=uv\gamma$.
  2. (ii)
    $D$ is ample if and only if u>0,v>0u>0,\,v>0.
  3. (iii)
    If $D$ is ample, then $h^{0}(\mathcal{O}_{S}(D))=\chi(\mathcal{O}_{S}(D))$.
  4. (iv)
    If $h^{0}(\mathcal{O}_{S}(D))\neq 0$, then $u\geq 0,\,v\geq 0$.
5.2.2. Proof of Proposition 5.1 in the bielliptic case
By Lemma 5.6 (i)–(iii), the numerical class of $\beta$ in $\operatorname{Num}(S)$ is of the form $a\lambda A+b\mu B$ with $a,b\in\mathbb{Z}_{>0}$, and
$\dim|d\beta|=\chi({\mathcal{O}}_{S}(d\beta))-1=d^{2}ab\lambda\mu\gamma-1.$
Any non-integral curve in $|d\beta|$ is in the image of a morphism from a scheme $P$ to $|d\beta|$, where $P$ parametrizes the triples $(C,D,L)$ satisfying $L\in\operatorname{Pic}^{0}(S)$, $C\in|C_{1}+L|$, and $D\in|C_{2}-L|$ for two nonzero effective divisors $C_{1}$, $C_{2}$ such that $C_{1}+C_{2}=d\beta$. By Lemma 4.9 (or Lemma 5.6), the images of finitely many such morphisms cover $|d\beta|\setminus|d\beta|^{\mathrm{int}}$. Thus, it suffices to estimate all possible values of
$\dim|d\beta|-\dim|C_{1}|\times|C_{2}|\times\operatorname{Pic}^{0}(S)=\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1.$
Suppose the numerical class of $C_{i}$ ($i=1,2$) is $a_{i}\lambda A+b_{i}\mu B$ ($a_{i},b_{i}\in\mathbb{Z}$). Then
$a_{1}+a_{2}=da,\quad b_{1}+b_{2}=db.$
It follows from Lemma 5.6 (iv) that $a_{i}\geq 0$ and $b_{i}\geq 0$.
Case 1: $a_{i}>0$ and $b_{i}>0$ for $i=1,2$. By Lemma 5.6 (ii) and (iii),
$\dim|C_{1}|+\dim|C_{2}|=\chi({\mathcal{O}}_{S}(C_{1}))+\chi({\mathcal{O}}_{S}(C_{2}))-2.$
Therefore, by the proof of (5.18), we have
(5.25) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1=C_{1}\cdot C_{2}\geq d\sqrt{2\beta^{2}}-2.$
Case 2: At least one of $a_{1},a_{2},b_{1},b_{2}$ is zero. We may assume $a_{1}=0$. Then $b_{1}>0$ since otherwise $C_{1}$ would be numerically trivial. By Lemma 5.6 (i)–(iii),
(5.26) $\dim|C_{1}|\leq\dim|A+C_{1}|=b_{1}\mu\gamma-1.$
If $b_{2}>0$, then it follows from Lemma 5.6 (i)–(iii), (5.26), and $b_{1}>0$ that
(5.27) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $=C_{1}\cdot C_{2}-\dim|C_{1}|-1$
$\geq dab_{1}\lambda\mu\gamma-(b_{1}\mu\gamma-1)-1$
$\geq(da\lambda-1)\mu\gamma\quad(\text{if }da\lambda\geq 1).$
If $b_{2}=0$, then as in (5.26), we have
$\dim|C_{2}|\leq\dim|C_{2}+B|=da\lambda\gamma-1,$
which, together with (5.26), implies
(5.28) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $\geq d^{2}ab\lambda\mu\gamma-1-(db\mu\gamma-1)-(da\lambda\gamma-1)-1$
$=d^{2}ab\lambda\mu\gamma-db\mu\gamma-da\lambda\gamma.$
Combining (5.25), (5.27), and (5.28) completes the proof. ∎
5.3. The case of K3 or Abelian surfaces
Now let $S$ be a K3 or an Abelian surface. The proof of Proposition 5.1 in this case is very similar to the proofs for Enriques and bielliptic surfaces (given in §5.1.2 and §5.2.2, respectively), so we keep it brief. Since the Kawamata–Viehweg vanishing theorem shows that Lemma 5.2 (i) is also valid for K3 and Abelian surfaces, we only need the following lemma (see [19, Theorem] for the K3 case, and [17, Proposition 2.3, Corollary 2.6] for the Abelian case).
Lemma 5.7.
Let $D$ be a nonzero effective nef divisor on a K3 or an Abelian surface $S$ with $D^{2}=0$. Then $D$ is numerically equivalent to $mE$ for some $m\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E$ with $E^{2}=0$. Moreover,
[…]
5.4. Some related stabilization questions
We collect some stabilization questions for moduli spaces of one-dimensional sheaves. Now $S$ is not necessarily a minimal surface of Kodaira dimension 0. For simplicity, we may assume $h^{1}({\mathcal{O}}_{S})=0$. We begin by providing a refined version of Conjecture 1.3, stated heuristically as follows.
Conjecture 5.8.
For fixed $i,j\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$, the perverse Hodge number $n_{h}^{i,j}$ of $h:M_{\beta,\chi}\rightarrow|\beta|$ stabilizes as $\beta$ becomes sufficiently positive. Furthermore, $n_{h}^{i,j}$ stabilizes to $H(q,t)]^{i,j}$ (defined as in Theorem 4.5) if $M_{\beta,\chi}$ is asymptotically irreducible.
In [36], the approach to Conjecture 1.3 for del Pezzo surfaces is to prove Conjecture 5.8. When $S$ is a K3 or an Abelian surface and $M_{\beta,\chi}$ is smooth, Shen–Yin [40] proved that the perverse Hodge numbers of $M_{\beta,\chi}$ are equal to its Hodge numbers using hyper-Kähler geometry and thus the stabilization of the perverse Hodge numbers follows from the closed formula of Hodge numbers of $S^{[n]}$ [12, Theorem 2.3.14]. Our Theorem 1.6 also shows the stabilization of perverse Hodge numbers when $S$ is an Enriques surface and $M_{\beta,\chi}$ is smooth.
Question 5.9.
For fixed $i,j\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$, does the intersection Hodge number $\dim_{\mathbb{Q}}I\!H^{i,j}(M_{\beta,\chi})$ stabilize as $\beta$ becomes sufficiently positive?
It seems that our method in this paper does not work for intersection Hodge numbers in the cases of Enriques and bielliptic surfaces. In [29], Mauri and Migliorini showed that Question 5.9 has a positive answer for K3 surfaces, based on the previously known case where $\chi=1$ and $S$ is a K3 surface.
The stabilization problems for moduli spaces of one-dimensional sheaves on $S$ can also be studied from a motivic perspective. Let $K_{0}(\mathrm{Var})$ be the Grothendieck ring of varieties, which is the free Abelian group generated by (the isomorphism classes of) varieties $[X]$ modulo the scissor relations
$[X]=[Y]+[Z],$
where $Y$ and $Z$ are disjoint locally closed subvarieties of $X$ with $X=Y\cup Z$. The product in $K_{0}(\mathrm{Var})$ is defined by taking the fiber product. Let $\mathbb{L}\in K_{0}(\mathrm{Var})$ be the class of the affine line $\mathbb{A}^{1}$. Denote by $\widehat{K}_{0}(\mathrm{Var})$ the completion of $K_{0}(\mathrm{Var})[\mathbb{L}^{-1}]$ with respect to the dimension filtration (see [8, p. 17] for more details).
Question 5.10.
For fixed $\chi\in\mathbb{Z}$, does the class $M_{\beta,\chi}]$ stabilize in $\widehat{K}_{0}(\mathrm{Var})$ as $\beta$ becomes sufficiently positive?
Coskun and Woolf [8] proved that when $S$ is a rational surface and $K_{S}\cdot H<0$, the class of the moduli space of $H$-stable torsion-free sheaves with fixed Chern character on $S$ stabilizes in $\widehat{K}_{0}(\mathrm{Var})$ as the determinant of sheaves tends to $\infty$. A positive answer to Question 5.10 would establish an analogous result in the rank 0 case.
A natural further question is whether there is a geometric interpretation for the stabilization of Betti numbers. One potential approach is to consider the tautological classes in $H^{*}(M_{\beta,\chi})$, i.e., the Künneth components of Chern classes of the universal sheaf (assuming its existence) on $M_{\beta,\chi}\times S$. Let $RH^{*}(M_{\beta,\chi})\subset H^{*}(M_{\beta,\chi})$ be the subring generated by the tautological classes. Although $RH^{*}(M_{\beta,\chi})\neq H^{*}(M_{\beta,\chi})$ in general (cf. [37, Theorem 5.10]), we expect the stabilization of Betti numbers of $M_{\beta,\chi}$ to be explained by the stabilization of $RH^{*}(M_{\beta,\chi})$. When $S$ is a del Pezzo surface and $M_{\beta,\chi}$ is smooth, one has $RH^{*}(M_{\beta,\chi})=H^{*}(M_{\beta,\chi})$ and the stable Betti numbers are determined by the combinatorics of tautological classes (see [36, Corollary 1.8]). If our expectation is true, the stabilization phenomenon in the del Pezzo case fits into a more general picture.
As for the intersection cohomology, the $k$-th stable intersection Betti number calculated in this paper also coincides with $b_{k}^{\infty}$. In what follows, we assume that $M_{\beta,\chi}$ is pure dimensional. Note that there is a homomorphism of graded vector spaces (e.g., [30, Exercise 6.7.1, Remark 6.7.2])
$\alpha:H^{*}(M_{\beta,\chi})\rightarrow I\!H^{*}(M_{\beta,\chi}).$
Thus we can define $RI\!H^{\ast}(M_{\beta,\chi})$ as the image of $RH^{\ast}(M_{\beta,\chi})$ under the map $\alpha$. Given our results, we also expect the stabilization of intersection Betti numbers to be explained by the stabilization of ${RI\!H}^{\ast}(M_{\beta,\chi})$, and therefore we ask the following question.
Question 5.11.
Is it true that $\dim_{\mathbb{Q}}{RI\!H}^{k}(M_{\beta,\chi})=b_{k}^{\infty}$ if $\beta$ is sufficiently positive?
Another direction is to explore the tautological relations in $RH^{\ast}(M_{\beta,\chi})$. When $S=\mathbb{P}^{2}$, the tautological relations are studied in [20], leading to the formulation of a strengthened version of the “P=C” conjecture proposed in [21]. For other surfaces, however, the problem remains unexplored.
More generally, one may consider similar questions for moduli spaces of one-dimensional sheaves on a projective surface with mild singularities. In this setting, we continue to use the notation $M_{\beta,\chi}$.
Question 5.12.
Fix $\chi\in\mathbb{Z}$ and let $\beta$ be a divisor on a simply connected projective surface with ADE singularities. For fixed $k\in\mathbb{Z}_{\geq 0}$, does the intersection Betti number $\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})$ stabilize as $\beta$ becomes sufficiently positive? Similarly, for fixed $i,j\in\mathbb{Z}_{\geq 0}$, does the intersection Hodge number $\dim_{\mathbb{Q}}I\!H^{i,j}(M_{\beta,\chi})$ stabilize?
Toward answering Question 5.12, we expect our strategy in this paper to work in some cases after suitable modifications, but there are some difficulties. For example, the relative Hilbert schemes of points over the linear system $|\beta|$ may not be equidimensional as the curves in $|\beta|$ may not be locally planar. Moreover, the Betti numbers (or intersection Betti numbers) of Hilbert schemes of points on a singular surface are still unknown.
Acknowledgments
The authors would like to thank Mirko Mauri, Weite Pi, and Junliang Shen for helpful discussions. FZ is grateful to her advisor Jun Li for his constant support, to János Kollár and Claire Voisin for inspiring discussions on the deformation theory of curve singularities, to Georg Oberdieck for suggesting considering the case of bielliptic surfaces, and to Songsong Huang for useful conversations.
FS is partially supported by the Fundamental Research Funds for the Central Universities, NSFC (No. 12025106) and Shaanxi NSF (No. 2025JC-QYCX-002). FZ is supported by the NSFC grants (No. 12121001 and No. 12425105) and LMNS.